Given that sin(theta)=3/5. Find sec(pi/2-theta).
Answer by Alan3354(69443)

Given that sin(theta)=3/5. Find sec(pi/2-theta).
-----
sin(t)=3/5
cos(90 - sin(t)) = 3/5
sec(90 - sin(t)) = 1/(cos) = 5/3